what is the radius in (mm) of a spherical water droplet with a volum of 0.088 cmm3
Answer #1

Volume of a sphene = 4 T43 > Radius. A = 3.147 V=0.088 cm3 0.088= 4 x 3.147 X 43 r² = 0.088* 3 p = 0·02097 cm3 4x 3-147 M = 4